# Better Fair than Sorry: Adversarial Missing Sensitive Value Imputation for Fair GNNs

### v2

@@ 3.3 Evaluation
Accuracy Evaluation: As with many real-world datasets in fairness literature, our datasets are imbalanced. To evaluate their accuracy, we apply average precision (AVPR) and the F1 score.
Fairness Evaluation: We apply two fairness metrics that will be defined based on a sensitive attribute $s$. Let $s=0$ be the protected group and $s=1$ be the non-protected group. The Demographic Parity (DP) requires the prediction to be independent of the sensitive attribute Kamiran and Calders (2009). We evaluate demographic parity as-
$\Delta DP(\hat{y},s)=|p(\hat{y}=1|s=1)-p(\hat{y}=1|s=0)|$ (9)
Equality of Opportunity (EQOP) calculates the difference between true positive rates between protected and non-protected groups Hardt et al. (2016). We evaluate Equality of Opportunity as-
$\begin{split}\Delta EQOP(\hat{y},s,y)=&|p(\hat{y}=1|s=1,y=1)\\ &-p(\hat{y}=1|s=0,y=1)|.\end{split}$
$\Delta DP$ and $\Delta EQOP$ calculate bias. We can get fairness by subtracting bias from $1$.
3.4 Hyperparameter settings
Each GNN has two layers and a dropout probability of $0.5$. Unless otherwise stated, $\mathcal{V}_{S}$ was set to $40\%$ of training data for all experiments. $\mathcal{V}_{L}$ is $300,400,1400,100$, and $800$ for the Simulation, German, Credit, NBA and Bail datasets, respectively. We adjust the values of $\alpha$ and $\beta$ for each dataset using cross-validation based on average precision. For all baselines, we choose the respective regularizers using cross-validation based on the same metrics. Every model, except RNF, combines the training of several networks with various convergence rates. We stop training RNF based on early stopping to eliminate overfitting. We train the rest of the models for certain epochs and choose the model with the best AVPR. All the results shown in the following section are the average of $10$ runs. We use GCN as the GNN architecture for all experiments unless mentioned otherwise.
@@ 3.5 Results and Analysis
Figure 3 shows the F1 score, Demographic Parity, and Equality of Opportunity of different methods while varying their parameter that controls the accuracy vs. fairness trade-off. We have generated the simulation graph using a Stochastic Block Model with different assortativity coefficients—i.e. the extent in which links exist within clusters compared with across clusters. Learning missing sensitive values and node labels with a low assortativity coefficient is hard and graphs with assortativity $0.17$ and $0.37$ represent the scenario described in Corollary 1, therefore, the JS divergence tends to be small. Figure 3 shows that for assortativity $0.17$ and $0.37$, FairGNN and Debias fail to learn the node labels in multiple cases. As we increase the assortativity coefficient, all benchmark models can learn the node labels. BFtS achieves a better fairness vs. accuracy trade-off in all cases.
Table 1 shows the average results for Bail, Credit, German, and NBA datasets. In terms of accuracy and fairness, BFtS outperforms all baseline approaches for the Bail dataset. Bail dataset is highly imbalanced with $7104$ positive samples amongst $18876$. Therefore, LDAM loss of the GNN classifier results in better AVPR and F1 than the Vanilla model. We also outperform competitors on the Credit dataset in terms of fairness with a slight AVPR loss. Compared to all other baselines, our bias is lower for the German dataset. To achieve this fairness, however, we sacrifice some AVPR. Although we lose some AVPR in the German and Credit datasets, we win in F1 and the AVPR loss is negligible when compared to the fairness benefit.
We demonstrate an accuracy and fairness trade-off in Figure 4. For BFtS, we vary $\alpha$, and $\beta$ to control the trade-off between accuracy and fairness. For other models, we vary their respective regularizers. On the x-axis of Figure 4, we have the F1 score. We have $1-\Delta DP$, and $1-\Delta EQOP$ on the y-axis of rows 1 and 2, respectively. Therefore, the top right corner of the plot represents higher fairness and accuracy. Our model achieves better fairness and similar accuracy for the Bail, NBA, German, and Simulation $(assortativity=0.67)$ datasets. For the Credit dataset, BFtS achieves competitive results against RNF and FairGNN.
@@
Figure 5 shows Kernel PCA node representations produced by FairGNN and BFtS using the Bail dataset. We show the samples with missing ”Race” with a navy dot in the middle. The figure shows samples predicted as positive class and negative class in green and red, respectively. The representations generated by FairGNN are noticeably more biased than the ones generated by our approach. We can see a higher number of ”Race = Black” samples that had the sensitive values missing in training are predicted as positive by BFtS than that of FairGNN. BFtS spreads nodes with missing values more uniformly over the space. The missing value imputation of FairGNN underestimated the bias in the dataset and therefore, the model could not overcome the bias in the predictions.
@@ 4 Conclusion
We present BFtS, a 3-player adversarial learning framework for the imputation of missing sensitive attributes that produces challenging values for graph-based fairness. Our theoretical analysis showed that our approach is more robust to convergence challenges during adversarial training than existing alternatives. We evaluate the BFtS architecture using synthetic and real datasets. Our experimental results provide strong evidence that our approach achieves a better fairness $\times$ accuracy trade-off than the baselines.

### v3

@@ 4.2 Evaluation
Accuracy Evaluation: As with many datasets in the fairness literature, our datasets are imbalanced. To evaluate their accuracy, we apply average precision (AVPR) and the F1 score.
Fairness Evaluation: We apply two bias metrics that will be defined based on a sensitive attribute $s$. Let $s=0$ be the protected group and $s=1$ be the non-protected group. The Demographic Parity (DP) requires the prediction to be independent of the sensitive attribute Kamiran and Calders (2009). We evaluate demographic parity as:
$\Delta DP(\hat{y},s)=|p(\hat{y}=1|s=1)-p(\hat{y}=1|s=0)|.$ (8)
Equality of Opportunity (EQOP) calculates the difference between true positive rates between protected and non-protected groups Hardt et al. (2016) as follows:
$\begin{split}\Delta EQOP(\hat{y},s,y)=&|p(\hat{y}=1|s=1,y=1)\\ &-p(\hat{y}=1|s=0,y=1)|.\end{split}$ (9)
@@ 4.3 Results and Analysis
Figure 4 shows the F1 score, $1-\Delta DP$, and $1-\Delta EQOP$ for different methods while varying their hyperparameter. We have generated the simulation graph using a stochastic block model with different assortativity coefficients, i.e., the extent to which links exist within clusters compared with across clusters. Learning missing sensitive values and node labels with a low assortativity coefficient is hard and graphs with assortativity $0.17$ and $0.37$ represent the scenario described in Corollary 1, therefore, the JS divergence tends to be small for independent imputation, which may result in a lack of convergence for the baselines. Figure 4 shows that for assortativity $0.17$ and $0.37$, FairGNN and Debias fail to predict node labels in multiple cases. As we increase the assortativity, all baselines can predict the labels, but BFtS still achieves a better fairness vs. accuracy trade-off in all cases.
@@
Figure 6: 2-D kernel PCA node representations for the Bail dataset generated by the FairGNN and BFtS (our approach). We use different markers for nodes depending on their predicted class, sensitive attribute (race), and whether the sensitive attribute is missing. The results show that representations for nodes with missing values are more biased for FairGNN where they are concentrated in a negative class region. On the other hand, nodes with missing values are better spread over the space for BFtS representations. In BFtS, there are more ”Race = Black” nodes than in FairGNN that have missing sensitive values in training but are predicted to be positive.
We demonstrate the accuracy vs. fairness trade-off for real datasets in Figure 5. For BFtS, we vary $\alpha$ and $\beta$ to generate different values of fairness and accuracy. For other models, we vary their respective regularizers. The top right corner of the plot represents high fairness and accuracy. Our model achieves better fairness and similar accuracy to the best baseline for the Bail, NBA, German, pokec-n and pokec-z. For all datasets, BFtS achieves a better fairness-accuracy tradeoff.
Figure 6 shows kernel PCA node representations produced by FairGNN and BFtS using the Bail dataset. We show the samples with missing ”Race” with a navy dot in the middle and samples predicted as positive class and negative class in green and red, respectively. The representations generated by FairGNN are noticeably more biased than the ones generated by our approach. A higher number of ”Race = Black” samples with missing sensitive values in training are predicted as positive by BFtS than for FairGNN. Moreover, BFtS spreads nodes with missing values more uniformly over the space. This illustrates how the missing value imputation of FairGNN underestimates the bias in the training data and, therefore, the model could not overcome the bias in the predictions.
Table 1 shows the accuracy and fairness results for Bail, Credit, German, Pokec-z, Pokek-n and NBA without any sensitive attribute information. Amongst the baselines, only RNF works in this setting. The results show that BFtS outperforms RNF for the Bail, German, and Credit. Our approach also outperforms RNF in terms of fairness on Pokec-z,Pokec-n, and NBA with similar AVPR and F1. As expected, we note that when we have $30\%$ of the sensitive attributes available (Figure 5), we obtain higher fairness than when we have no sensitive attribute information (Table 1).
Table 2 shows the imputation accuracy of the methods. We exclude results for Debias, which is based only on the sensitive information available, and for FairVGNN and FairSIN, as they are identical to FairGNN (same imputation method). The results show that BFtS imputation outperforms independent imputation methods. BFtS worst-case imputation based on the LDAM loss outperforms the alternatives in terms of accuracy. This is due to the adversarial missingness process where low-degree nodes are selected to have missing values. Independent imputation methods are less effective in this adversarial setting.
German Credit Bail NBA pokec-z pokec-n
FairGNN 0.70 0.90 0.56 0.78 0.81 0.84
RNF 0.59 0.73 0.53 0.64 0.65 0.68
BFtS 0.72 0.94 0.63 0.78 0.83 0.87
Table 2: Accuracy of missing sensitive imputation.
@@ 5 Conclusion
We present BFtS, a 3-player adversarial learning framework for the imputation of missing sensitive attributes that produce challenging values for graph-based fairness. Our theoretical analysis showed that our approach is more robust to convergence challenges during adversarial training than existing alternatives. We evaluate BFtS using synthetic and real datasets. Our experimental results provide strong evidence that our approach achieves a better fairness $\times$ accuracy trade-off than the baselines.